OpenStudy (anonymous):

Use VSEPR theory to predict bond angles in the following covalent bond molecules. a. methane b. ammonia c. water

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@amistre64

OpenStudy (anonymous):

@hartnn

OpenStudy (anonymous):

@hba @mark_o.

OpenStudy (anonymous):

methane = CH4 Shape = (4 + 4)/2 = 4 tetrahedral angle 109.5

OpenStudy (anonymous):

ammonia = NH3 shape = (5 +3)/2 = 4 but it has 1 lone pair (repulsion) angle 107.5

OpenStudy (anonymous):

what is the "bond angle" part

OpenStudy (anonymous):

water H20 (6 + 2)/2 it has 2 Lp So bond angle is 104.5

OpenStudy (anonymous):

hope this helps

OpenStudy (anonymous):

how do you find the bond angle?

OpenStudy (anonymous):

@Yahoo! how do you find the bond angle? i dont understand....

OpenStudy (anonymous):

@sirm3d can you help me?

OpenStudy (sirm3d):

the angles between bonds were computed using solid geometry, and the results were provided as facts in textbooks.